If a*b = a²-b² find (i) 8*20 (II) (-4*6) -7If a*b = a÷b + b÷a evaluate 3*4

Grade:6

1 Answers

Arun
6 years ago
in 1st part,
we have to simply put a =8 & b = 20
now 8*20 = 82 –202 = 64 – 200 =  – 136
 
in 2nd part also , we have to put a = 3 & b = 4
now 3 * 4 = (3/4) + (4/3) = (9 + 16) / 12 = 25/12
 
hope it helps